OpenStudy (anonymous):

What is an equation of a parabola with the given vertex and focus? vertex: (–2, 5); focus: (–2, 6) (3 points)

OpenStudy (anonymous):

y-k = (-1/(4p))(x-h)² vertex (h,k) focus (h,k-p) h=-2 k=5 5-p=6 where p=-1 y-5=-1/(4*(-1))(x+2)^2 y-5=(1/4)(x+2)^2